Statistical Analysis: Correlation and Descriptive Statistics

Write results in APA Format.

Sleep blks Sleep Qual # correct
36 5 4
48 7 13
43 7 12
32 7 12
38 7 11
44 5 11
58 5 11
54 6 9
45 5 6
27 7 12
54 4 11
38 4 8
27 5 10
42 5 6
38 6 10
48 5 13
54 7 11
45 8 10
60 6 5
56 7 8
48 6 10
50 7 12
60 7 6
45 5 5
43 5 5
43 6 8
33 5 7
51 8 12
37 10 5
51 6 8
54 7 13
45 6 11
48 7 10
63 7 9
54 7 8
46 6 10
30 9 10
48 10 7
Answer #1

here number of observations n=38 and correlation matrix

Sleep blks Sleep Qual # correct
Sleep blks 1
Sleep Qual 0.009187085 1
# correct -0.008826377 0.10506868 1

here we want to test the correlation coefficient Sleep blks and Sleep Qual.

r=0.0092 ( four decimal place approximation)

Null hypothesis H0:ρ=0 and

Alternate hypothesis Ha: ρ ≠ 0

and we use t-test to test the above pair of hypothesis and

statistic t =r/sqrt[(1—r2)/(n—2)] =0.0092/sqrt((1-0.0092*0.0092)/(38-2))=0.0552 with n-2=38-2=36 degree of freedom(df)

the typical two tailed critical t(0.05,36)=2.0281 is more than calculated t=0.0552 so we fail to reject H0 and conclude that there is 0 correlation between sleep blks and sleep qual.